# Perturb-and-Project: Differentially Private Similarities and Marginals

### v2

@@ 4 Techniques
$\operatorname*{\mathbb{E}}_{\mathbf{W}\sim N(0,\mathrm{Id}_{n})^{n\times n}}\left[\sup_{X\in\mathcal{S}}\langle X,\mathbf{W}\rangle\right]\leqslant\operatorname*{\mathbb{E}}\left[\left\lVert X\right\rVert_{\textnormal{nuc}}\left\lVert\mathbf{W}\right\rVert\right]\leqslant n\cdot 2\sqrt{n}\cdot\sigma\,,$
@@
$\left\{X\in(\mathbb{R}^{n})^{\otimes k}\;\middle|\;\sum_{i\in[m]}v_{i}^{\otimes k}\,,\left\lVert X\right\rVert_{\textnormal{max}}\leqslant 1\right\}\,.$ (3)
@@
[…]
@@ 5 Guarantees of the perturb-and-project paradigm for differential privacy
$\left\lVert A-\Pi(A)\right\rVert^{2}_{2}$
$\geqslant\left\lVert A-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}$
$+\langle\nabla\left\lVert A-\Pi(A+\mathbf{W})\right\rVert^{2}_{2},\Pi(A)-\Pi(A+\mathbf{W})\rangle$
$+\frac{1}{2}\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$
@@
$\left\lVert A+\mathbf{W}-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}$
$\geqslant\left\lVert A+\mathbf{W}-\Pi(A)\right\rVert^{2}_{2}$
$+\langle\nabla\left\lVert A+\mathbf{W}-\Pi(A)\right\rVert^{2}_{2},\Pi(A+\mathbf{W})-\Pi(A)\rangle$
$+\frac{1}{2}\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}\,.$
@@
$\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$
$\leqslant\langle\nabla\left\lVert A-\Pi(A+\mathbf{W})\right\rVert^{2}_{2},\Pi(A)-\Pi(A+\mathbf{W})\rangle$
$+\langle\nabla\left\lVert A+\mathbf{W}-\Pi(A)\right\rVert^{2}_{2},\Pi(A+\mathbf{W})-\Pi(A)\rangle$
$=-2\left\lVert\Pi(A)-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}$
$+2\langle\mathbf{W},\Pi(A+\mathbf{W})-\Pi(A)\rangle\,.$
@@
$\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$
$\leqslant\frac{2}{3}\operatorname*{\mathbb{E}}\langle\mathbf{W},\Pi(A+\mathbf{W})-\Pi(A)\rangle$
$\leqslant\frac{4}{3}\operatorname*{\mathbb{E}}\mathcal{G}(\mathcal{S})\,.$
@@ 5.1 Practical perturb-and-project algorithms
$\operatorname*{\mathbb{E}}\left\lVert\hat{\mathbf{\Pi}}_{t}-\Pi_{\mathcal{S}}(A)\right\rVert^{2}_{2}\leqslant$ $\frac{16\sqrt{\log(2/\delta)}}{3\varepsilon}\cdot\Delta$
$\cdot\left(\mathcal{G}(\mathcal{S})+2^{-t}\left\lVert A\right\rVert^{2}_{2}+n^{-\Omega(1)}\right)\,.$
@@
$\operatorname*{\mathbb{E}}_{\mathbf{W}}$ $\left\lVert\Pi_{t}(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$
$\leqslant 2\operatorname*{\mathbb{E}}_{\mathbf{W}}\left\lVert\Pi_{t}(A+\mathbf{W})-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}$
$+2\operatorname*{\mathbb{E}}_{\mathbf{W}}\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$
@@
$\operatorname*{\mathbb{E}}_{\mathbf{W}}$ $\left\lVert\Pi_{t}(A+\mathbf{W})-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}$
$\leqslant\operatorname*{\mathbb{E}}_{\mathbf{W}}c^{t}\left\lVert A+\mathbf{W}-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}$
@@ Acknowledgments
Impact statement
This paper presents work whose goal is to advance the field of differentially private machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v3

@@ 4 Techniques
$\operatorname*{\mathbb{E}}_{\mathbf{W}\sim N(0,\mathrm{Id}_{n})^{n\times n}}\left[\sup_{X\in\mathcal{S}}\langle X,\mathbf{W}\rangle\right]$ $\leqslant\operatorname*{\mathbb{E}}\left[\left\lVert X\right\rVert_{\textnormal{nuc}}\left\lVert\mathbf{W}\right\rVert\right]$
$\leqslant n\cdot 2\sqrt{n}\cdot\sigma\,,$
@@
$\left\{X\in(\mathbb{R}^{n})^{\otimes k}\;\middle|\;\sum_{i\in[m]}v_{i}^{\otimes k}=X\,,\left\lVert X\right\rVert_{\textnormal{max}}\leqslant 1\right\}\,.$ (3)
@@
$\left\{X\in(\mathbb{R}^{n})^{\otimes k}\,|\,\exists\textnormal{ deg-$O(k)$ }\mu\textnormal{ over }\mathcal{Q}\textnormal{ s.t. }\tilde{\operatorname*{\mathbb{E}}}_{\mu}[\sum_{i\in[m]}v_{i}^{\otimes k}]=X\right\}\,,$
@@ 5 Guarantees of the perturb-and-project paradigm for differential privacy
$\left\lVert A-\Pi(A)\right\rVert^{2}_{2}$ $\geqslant\left\lVert A-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}+\langle\nabla\left\lVert A-\Pi(A+\mathbf{W})\right\rVert^{2}_{2},\Pi(A)-\Pi(A+\mathbf{W})\rangle+\frac{1}{2}\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$
@@
$\left\lVert A+\mathbf{W}-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}$ $\geqslant\left\lVert A+\mathbf{W}-\Pi(A)\right\rVert^{2}_{2}$
$+\langle\nabla\left\lVert A+\mathbf{W}-\Pi(A)\right\rVert^{2}_{2},\Pi(A+\mathbf{W})-\Pi(A)\rangle+\frac{1}{2}\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}\,.$
@@
$\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$ $\leqslant\langle\nabla\left\lVert A-\Pi(A+\mathbf{W})\right\rVert^{2}_{2},\Pi(A)-\Pi(A+\mathbf{W})\rangle+\langle\nabla\left\lVert A+\mathbf{W}-\Pi(A)\right\rVert^{2}_{2},\Pi(A+\mathbf{W})-\Pi(A)\rangle$
$=-2\left\lVert\Pi(A)-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}+2\langle\mathbf{W},\Pi(A+\mathbf{W})-\Pi(A)\rangle\,.$
@@
$\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$ $\leqslant\frac{2}{3}\operatorname*{\mathbb{E}}\langle\mathbf{W},\Pi(A+\mathbf{W})-\Pi(A)\rangle\leqslant\frac{4}{3}\operatorname*{\mathbb{E}}\mathcal{G}(\mathcal{S})\,.$
@@ 5.1 Practical perturb-and-project algorithms
$\operatorname*{\mathbb{E}}\left\lVert\hat{\mathbf{\Pi}}_{t}-\Pi_{\mathcal{S}}(A)\right\rVert^{2}_{2}\leqslant$ $\frac{16\sqrt{\log(2/\delta)}}{3\varepsilon}\cdot\Delta\cdot\left(\mathcal{G}(\mathcal{S})+2^{-t}\left\lVert A\right\rVert^{2}_{2}+n^{-\Omega(1)}\right)\,.$
@@
$\operatorname*{\mathbb{E}}_{\mathbf{W}}\left\lVert\Pi_{t}(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$ $\leqslant 2\operatorname*{\mathbb{E}}_{\mathbf{W}}\left\lVert\Pi_{t}(A+\mathbf{W})-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}+2\operatorname*{\mathbb{E}}_{\mathbf{W}}\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$
@@
$\operatorname*{\mathbb{E}}_{\mathbf{W}}\left\lVert\Pi_{t}(A+\mathbf{W})-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}$ $\leqslant\operatorname*{\mathbb{E}}_{\mathbf{W}}c^{t}\left\lVert A+\mathbf{W}-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}$